# LightGaussian: Unbounded 3D Gaussian Compression with 15x Reduction and 200+ FPS

### v4

@@ 6 Acknowledgements
[…]
9:#Distilling into Compact SHs.
[…]
@@
13:    $I_{t}\leftarrow\textsc{\text{Teacher}}(\mathcal{\hat{P}})$triangleright Teacher render
14:    $I_{s}\leftarrow\textsc{\text{Student}}(\mathcal{\hat{P}})$triangleright Student render
@@
16:    $\mathcal{\hat{G}}\leftarrow\textsc{Adam}(\nabla L)$triangleright Backprop & Step
@@
18:#VecTree Quantization.
[…]
21:$\text{XYZ}^{\prime}\leftarrow$ OctreeCompression(XYZ)
22:#Save Model.
23:Save optimized model $\mathcal{\hat{G}}$ to disk.
@@
2:# $\mathcal{P}$ Sample Camera poses from training.
@@
4:    $GS\leftarrow 0$triangleright Init Global Significance
5:    for all pixels $\ i\ \ \text{in renderFunc}(\mathcal{P})$ do
6:         $H\leftarrow$ GetHitCount($\mathcal{G},i$) $\triangleright$ Gaussian Hit count
7:         # Eq.5 in the main draft.
8:         $GS\leftarrow GS+H\cdot Opacity^{\intercal}\cdot$ VNorm($Scale$${}^{\intercal}$
9:    end for
10:    return $GS$
11:end function
A8 More Experiment Results
@@
The synthetic Blender dataset [43] includes eight photo-realistic synthetic objects with ground-truth controlled camera poses and rendered viewpoints (100 for training and 200 for testing). Similar to 3D-GS [32], we start training the model using random initialization. Consequently, we calculate the Global Significance of each Gaussian, work to reduce the SH redundancy, and apply the VecTree Compression (codebook size set at 8192) to the learned representation. Overall comparisons with previous methods are listed in Table A6, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
Table A6: Per-scene results on Synthetic-NeRF.
@@
We provide extra visualizations for 3D-GS [32], LightGaussian (ours), and VQ-DVGO [38], accompanied by the corresponding residual maps from the ground truth. As evidenced in Figure A8 and Figure A9, LightGaussian outperforms VQ-DVGO, which utilizes NeRF as a basic representation. Furthermore, LightGaussian achieves a comparable rendering quality to 3D-GS [32], demonstrating the effectiveness of our proposed compact representation.
@@
Tables A7, A8, and A9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 727MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
Figure A8: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [38]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Figure A9: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [38]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Table A7: Quantitative Comparison (PSNR) for Mip-NeRF360 and Tank & Temple Scenes.
@@
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 21.912 23.4947 20.661 27.594 23.624 23.420 24.669 23.221 18.927
INGP-Big 22.171 25.069 23.466 29.690 26.691 29.479 30.685 23.383 20.456
mip-NeRF360 24.305 26.875 26.175 31.467 29.447 31.989 33.397 24.912 19.523
VQ-DVGO 22.089 24.119 23.455 28.423 26.084 25.930 27.985 - -
3D-GS 25.122 27.294 26.783 31.687 29.114 31.618 32.307 24.778 21.449
Ours 24.960 26.735 26.701 31.271 28.113 30.402 31.014 24.561 21.095
@@
Table A8: Quantitative Comparison (SSIM) for Mip-NeRF360 and Tank & Temple Scenes.
@@
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.496 0.6063 0.523 0.8417 0.759 0.648 0.814 0.774 0.663
INGP-Big 0.512 0.701 0.594 0.871 0.817 0.858 0.906 0.800 0.689
mip-NeRF360 0.685 0.809 0.745 0.910 0.892 0.917 0.938 0.857 0.660
VQ-DVGO 0.473 0.613 0.564 0.860 0.777 0.763 0.842 - -
3D-GS 0.746 0.856 0.770 0.914 0.914 0.932 0.946 0.863 0.781
Ours 0.738 0.836 0.768 0.926 0.893 0.914 0.933 0.855 0.760
Table A9: Quantitative Comparison (LPIPS) for Mip-NeRF360 and Tank & Temple Scenes.
@@
bicycle garden stump room counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.506 0.3864 0.503 0.4186 0.441 0.447 0.398 0.335 0.422
INGP-Big 0.446 0.257 0.421 0.261 0.306 0.195 0.205 0.249 0.360
mip-NeRF360 0.305 0.171 0.265 0.213 0.207 0.128 0.179 0.159 0.354
VQ-DVGO 0.572 0.404 0.424 0.219 0.244 0.219 0.191 - -
3D-GS 0.245 0.122 0.242 0.200 0.186 0.118 0.184 0.175 0.262
Ours 0.265 0.155 0.261 0.220 0.218 0.147 0.204 0.188 0.296

### v5

@@ 5 More Technical Details
5:#Gaussian Pruning and Recovery: $\mathcal{G}\mapsto\mathcal{G^{\prime}}$.
@@
7:$\mathcal{G^{\prime}}\leftarrow\textsc{Prune}(\mathcal{G},\mathcal{GS})$triangleright Prune Least Significant Ones of $\mathcal{G}$, based on $\mathcal{GS}$
8:$\mathcal{G^{\prime}}\leftarrow\textsc{Recovery}(\mathcal{G^{\prime}},\mathcal{P})$$\triangleright$ Gaussian Recovery (Finetune $\mathcal{G^{\prime}}$ on $\mathcal{P}$)
9:#Distilling into Compact SHs: Reduce from 3-degree to 2-degree
10:$\mathcal{G^{\prime\prime}}\leftarrow\textsc{ReduceSH}(\mathcal{G^{\prime}})$triangleright Reduce the SH degree
@@
13:    $I_{t}\leftarrow\textsc{\text{Teacher}}(\mathcal{\hat{P}},\mathcal{G^{\prime}})$triangleright Teacher render
14:    $I_{s}\leftarrow\textsc{\text{Student}}(\mathcal{\hat{P}},\mathcal{G^{\prime\prime}})$triangleright Student render
@@
16:    $\mathcal{G^{\prime\prime}}\leftarrow\textsc{Adam}(\nabla L)$triangleright Backprop & Step
@@
18:#Vector Quantization.
19:$\mathcal{G^{\prime\prime\prime}}\leftarrow$ VQ($\mathcal{G^{\prime\prime}},\mathcal{GS}$) $\triangleright$ VQ based on Significance Score
20:$\mathcal{G^{\prime\prime\prime}}\leftarrow$ VQFineTune($\mathcal{G^{\prime\prime\prime}}$) $\triangleright$ VQ Finetuning
21:#Save Model.
22:Save optimized model $\mathcal{G^{\prime\prime\prime}}$ to disk.
@@
2:# $\mathcal{P}$ Sample cameras from training views.
@@
4:    ScoreList = list()
5:    for all Gaussian ${}_{i}\ \text{in}\ \mathcal{G}$ do
6:         $\text{score}\leftarrow 0$triangleright Init Global Significance
7:         $\text{count}\leftarrow 0$triangleright Init Hit Count Value
[…]
9:             count = GetHitCount(Gi, $r_{i}$) $\triangleright$ Gaussian Hit count
10:             # Score weighted by opacity/volume
[…]
12:         end for
13:         ScoreList.append(score)
14:    end for
15:    return ScoreList
16:end function
Post-Finetuning of Vector Quantization.
Applying vector quantization (VQ) based on the significance score facilitates a reduction in the required codebook size compared to the full model. Nonetheless, this method results in a decrease in accuracy as it groups the attributes into several clusters (codes). We observe that the SSIM decreases from 0.923 (#6) to 0.915 (#7), as illustrated in Table 2 of the main draft. This observation motivates us to joint optimize all other attributes and the codes post the VQ. Specifically, by assigning a unique code to each Gaussian, we proceed to implement differentiable finetuning using photometric loss, which is similar to the methodology employed in 3D-GS, and optimize all the attributes in conjunction with VQ. This approach permits precise adjustments to the codebook and the rest attributes, improving its alignment with the training images (#8, SSIM improves from 0.915 to 0.923).
6 Detailed Experiment Results
@@
The synthetic Blender dataset [45] includes eight photo-realistic synthetic objects with ground-truth controlled camera poses and rendered viewpoints (100 for training and 200 for testing). Similar to 3D-GS [32], we start training the model using random initialization. Consequently, we calculate the Global Significance of each Gaussian, work to reduce the SH redundancy, and apply the Vector Quantization (codebook size set at 8192) to the learned representation. Overall comparisons with previous methods are listed in Table 6, where we observe our LightGaussian markedly reduces the average storage size from 52.38MB to 7.89MB, while improving the FPS from 310 to 411 with only a slight rendering quality decrease.
Table 6: Per-scene results on Synthetic-NeRF.
@@
We provide extra visualizations for 3D-GS [32], LightGaussian (ours), and VQ-DVGO [40], accompanied by the corresponding residual maps from the ground truth. As evidenced in Figure 8 and Figure 9, LightGaussian outperforms VQ-DVGO, which utilizes NeRF as a basic representation. Furthermore, LightGaussian achieves a comparable rendering quality to 3D-GS [32], demonstrating the effectiveness of our proposed compact representation.
@@
Tables 7, 8, and 9 present the comprehensive error metrics compiled for our evaluation across all real-world scenes (Mip-NeRF360 and Tank and Temple datasets). Our method not only compresses the average model size from 724MB to 42MB, but also consistently demonstrates comparable metrics with 3D-GS on all scenes. LightGaussian additionally shows better rendering quality than Plenoxel, INGP, mip-NeRF360, and VQ-DVGO.
Figure 8: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Figure 9: Additional Visual Comparisons on the Mip-NeRF360 Datasets. We present the rendering results from 3D-GS [32], LightGaussian, and VQ-DVGO [40]. The corresponding residual maps highlighting the differences between the rendered images and ground truth (GT) images are also displayed.
Table 7: Quantitative Comparison (PSNR) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 21.912 20.097 23.4947 20.661 22.248 27.594 23.624 23.420 24.669 23.221 18.927
INGP-Big 22.171 20.652 25.069 23.466 22.373 29.690 26.691 29.479 30.685 23.383 20.456
mip-NeRF360 24.305 21.649 26.875 26.175 22.929 31.467 29.447 31.989 33.397 24.912 19.523
VQ-DVGO 22.092 18.934 24.127 23.428 28.502 21.440 26.035 25.459 27.990 - -
3D-GS 25.122 21.324 27.294 26.783 31.687 22.500 29.114 31.618 32.307 24.778 21.449
Ours 25.093 21.252 27.043 26.862 31.680 22.658 28.464 30.974 31.459 24.851 21.407
@@
Table 8: Quantitative Comparison (SSIM) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.496 0.431 0.606 0.523 0.509 0.842 0.759 0.648 0.814 0.774 0.663
INGP-Big 0.512 0.486 0.701 0.594 0.542 0.871 0.817 0.858 0.906 0.800 0.689
mip-NeRF360 0.685 0.584 0.809 0.631 0.745 0.910 0.892 0.917 0.938 0.857 0.660
VQ-DVGO 0.473 0.355 0.614 0.563 0.861 0.476 0.777 0.760 0.842 - -
3D-GS 0.746 0.574 0.856 0.770 0.926 0.630 0.914 0.932 0.946 0.863 0.781
Ours 0.745 0.566 0.845 0.775 0.923 0.627 0.900 0.921 0.938 0.862 0.772
Table 9: Quantitative Comparison (LPIPS) on Mip-NeRF360 and Tank & Temple Scenes. We re-train 3D-GS using their original code and configurations on our platform, to perform a fair comparison.
@@
bicycle flowers garden stump room treehill counter kitchen bonsai truck(T&T) train(T&T)
Plenoxels 0.506 0.521 0.3864 0.503 0.540 0.4186 0.441 0.447 0.398 0.335 0.422
INGP-Big 0.446 0.441 0.257 0.421 0.450 0.261 0.306 0.195 0.205 0.249 0.360
mip-NeRF360 0.305 0.346 0.171 0.265 0.347 0.213 0.207 0.128 0.179 0.159 0.354
VQ-DVGO 0.571 0.628 0.402 0.425 0.216 0.640 0.244 0.222 0.191 - -
3D-GS 0.245 0.381 0.122 0.242 0.200 0.366 0.186 0.118 0.184 0.175 0.262
Ours 0.257 0.392 0.143 0.251 0.212 0.378 0.211 0.140 0.199 0.179 0.283
7 Conclusion and Future Works
We present LightGaussian, a novel framework that converts the heavy point-based representation into a compact format for efficient novel view synthesis. For practical use, LightGaussian explores the use of 3D Gaussians for modeling large-scale scenes and finds an effective way to identify the least important Gaussians grown by densification. In pursuing a compact Spherical Harmonics format, distillation is enhanced by synthesizing pseudo-views to generate more data for knowledge transfer. A 3D Gaussian Vector Quantization post-processing effectively removes further redundancy by using Gaussian significance as an indicator. With these methods, the proposed representation reduces data redundancy by more than 15$\times$, further boosting the FPS to more than 200 FPS, with minimal rendering quality loss. Exploring the possibility of zero-shot compression across various 3D-GS-based frameworks is a valuable direction for future research.